OpenStudy (anonymous):

If y=2sqrt(x)-(1/2sqrt(x)), then what is the derivative of y with respect to x?

OpenStudy (amorfide):

square root means power of a half \[\sqrt{x}=x^{\frac{ 1 }{ 2 }}\]

OpenStudy (amorfide):

\[\frac{ 1 }{ 2\sqrt{x} }= \frac{ 1 }{ 2}\frac{ 1 }{ \sqrt{x} }\] \[\frac{ 1 }{ a}=a^{-1}\]

OpenStudy (amorfide):

so you end up re writing your equation to be \[y=2x^{\frac{ 1 }{ 2 }} - \frac{ x^{-\frac{ 1 }{ 2 }} }{ 2 }\]

OpenStudy (amorfide):

now you can differentiate this using \[\frac{ dy }{ dx }=nx^{n-1}\] where n is the power

OpenStudy (amorfide):

give it a go and I will correct you if you go wrong